OpenStudy (anonymous):

Solve. Please show the steps to solve the equation 4(x - 6) = -2(x - 3)

OpenStudy (saifoo.khan):

\[\frac42(x-6) = -(x-3)\] 2(x-6) = -x+3

OpenStudy (saifoo.khan):

2x - 12 = -x + 3 2x + x = 3 + 12 3x = 15 x = 15/3 x = Answer.

OpenStudy (anonymous):

4(x-6) = -2(x-3) Distribute. 4x-24 = -2x+6 Isolate the variable. 6x=30 x=5